# Robustifying Long-term Human-Robot Collaboration through a Hierarchical and Multimodal Framework

### v1

@@ V-C4 H4 - Adaptibility
We employ an online adaptation strategy for the trajectory prediction model to refine its predictions. We conducted experiments to evaluate the performance of an offline-trained prediction model across four different users. Accuracy was used to evaluate intention prediction, while mean-squared-error (MSE) was used to evaluate trajectory prediction. Table IV presents the results of predictions with and without online adaptation. The results clearly show that online adaptation improves the performance of both intention and pose prediction. The improvement in trajectory prediction is particularly significant, with a 24% reduction in MSE. This is expected, as different individuals may have varying pose and pose preferences. The results demonstrate that the proposed framework effectively adapts to different human behaviors.
@@
TABLE IV: Prediction results with and without online adaptation.
@@ V-C5 H5 - User Satisfaction
We conducted a user study to evaluate user satisfaction with the toy car assembly HRC task, collecting responses to evaluate the effectiveness of the proposed multimodal framework. Figure 9 participants’ responses to six questions across three different HRC frameworks: vision-based, audio-based, and the proposed multimodal framework. As shown, the percentage of participants who selected ”strongly agree” was significantly higher for the multimodal framework compared to the single-modal frameworks. For example in Figure 9(f), many users ”strongly agree” with the statement, ”I am satisfied with this framework,” when using the multimodal framework. Participants rated their agreement on a scale from 1 (strongly disagree) to 5 (strongly agree). The average satisfaction score for each framework across the six questions is summarized in section -B.
Figure 9: User Study
@@
In the user study, the multimodal framework received the highest satisfaction ratings from most participants, outperforming the single-modal frameworks. Specifically, user feedback highlighted that: 1) Users found it easier to interact with the robot using the proposed multimodal framework, indicating an improved user experience. 2) Users expressed a greater willingness to collaborate with the multimodal framework over an extended period, suggesting that it effectively reduces user workload. 3) Overall, users reported higher satisfaction when working with the multimodal framework. This user study demonstrates the effectiveness of the multimodal framework in creating a user-friendly HRC system for long-term HRC tasks.
VI CONCLUSIONS
In this work, we propose a hierarchical and multimodal framework for human-robot collaboration tasks. Vision observations from the environment are used to predict human intentions through hierarchical human pose detection and intention prediction, while audio observations are employed to recognize speech commands. The predicted intentions and recognized speech are then used to guide the robot’s motions according to the plan graph. Experiments and user studies demonstrate that the proposed framework significantly improves the efficiency of long-term HRC tasks and increases user satisfaction.

### v2

@@ V-C4 H4 - Adaptability
We employ an online adaptation strategy for the human physical action prediction model to refine its predictions. We conducted experiments to evaluate the performance of an offline-trained prediction model across four different users. Accuracy was used to evaluate intention prediction, while mean-squared-error (MSE) was used to evaluate trajectory prediction. Table. V presents the results of predictions with and without online adaptation. The results clearly show that online adaptation improves the performance of both action and trajectory prediction. The improvement in trajectory prediction is particularly significant, with a 24% reduction in MSE. This is expected, as different individuals may have varying pose and pose preferences. The results demonstrate that the proposed framework effectively adapts to different human behaviors.
@@
TABLE V: Prediction results with and without online adaptation.
@@ V-C5 H5 - User Satisfaction
Fig. 9 shows the user plan satisfaction rate collected from six groups across three subtasks. As can be seen from the figure, with hierarchical plan prediction and multimodal perception (”PP = 1, PM = 2”), we can achieve the largest user plan satisfaction rate. Besides, with hierarchical plan prediction (”PP = 1”), we can improve the user plan satisfaction rate by 47.5% than without hierarchical plan prediction (”PP = 0”). And with multimodal perception (”PM = 2”), we can improve the user plan satisfaction rate by 27.9% than without multimodal perception (”PM = 0” and ”PM = 1”). These results prove that both the hierarchical plan prediction method and the multimodal perception can significantly improve the user plan satisfaction rate.
Figure 9: The user plan satisfaction rate, as calculated based on 120 trials across all three sub-tasks for each group by comparing the recorded user plan sequences with the actual execution sequences.
Figure 10: Quantified results of the user survey, with scores calculated as weighted sums of responses. Each question was rated on a five-point scale: ”Strongly Disagree: 1”, ”Disagree: 2”, ”Neutral: 3”, ”Agree: 4”, and ”Strongly Agree: 5”.
To gain deeper insights into the user experience with our framework, we conducted a survey featuring six carefully designed questions: timely responsiveness (Q1), misunderstanding rates (Q2), ease of interaction (Q3), support for maintaining focus on the primary task (Q4), willingness to work with the framework for extended periods (Q5) and overall satisfaction (Q6). These questions evaluate the framework’s user-friendliness and overall effectiveness, providing valuable insights into its ability to enhance user flexibility and satisfaction.
@@
Fig. 10 shows quantized user satisfaction scores for six groups across six survey questions. Specifically, user feedback highlighted that: 1) Users found the framework highly responsive, accurate, and easy to interact with (Q1, Q2, Q3), reflecting an improved overall user experience. 2) Users reported greater focus on primary tasks rather than the collaboration process with the robot (Q4), indicating a higher user flexibility. 3) Users expressed a greater willingness to collaborate with the hierarchical and multimodal framework over an extended period (Q5), suggesting that it effectively reduces user workload. 4) Overall, users reported higher satisfaction when working with the hierarchical and multimodal framework (Q6). This user study demonstrates the effectiveness of the proposed hierarchical and multimodal framework in creating a user-friendly HRC system for long-term HRC tasks.
VI LIMITATIONS
While this work significantly showcases its great efficacy in accurate understanding of human-related behaviors, it does not address object-related understanding. However, an accurate understanding of real-time assemble structure would greatly contribute to better human plan prediction and task progree inference. Also, in this work, we assume no uncertainty regarding the location of objects. However, minimizing object-related uncertainties remains a substantial challenge for future work. Furthermore, future work could explore the use of powerful world models to model human behaviors for more flexible understanding of human actions and plans, though this may require a trade-off between flexibility and precision. Additionally, the task graph in this work requires manual design for specific scenarios; automating this process in future research would improve scalability to diverse tasks.
VII CONCLUSIONS
In this work, we identify the key challenges of long-term HRC tasks, present a structured task graph and propose a multimodal and hierarchical framework to address these challenges. We deploy our framework in a real-world long-term HRC assembly scenario and conducted extensive experiments and user studies. The result demonstrate that the proposed framework significantly improves the accuracy, robustness, efficiency, adaptability, and human satisfaction during the long-term HRC process, demonstrating the framework’s great potential to further long-term HRC applications.